# LAENeRF: Local Appearance Editing for Neural Radiance Fields

### v1

@@ 5.1 Photorealistic Appearance Editing
For the quantitative evaluation, we follow ICE-NeRF [23] and measure the Mean Squared Error (MSE) in the background of the selected region before and after recoloring. In Tab. 1, we present this metric for three scenes of the LLFF dataset [27]. The foreground region is described by a mask, which was provided to us by the authors of ICE-NeRF. We perform 7 different recolorings per scene and compare against PaletteNeRF with and without semantic guidance, whereas we also include the numbers from ICE-NeRF [^1] ICE-NeRFs’ implementation is not publicly available at the time of submission. to facilitate cross-method comparisons. LAENeRF outperforms previous methods for this metric, reducing error rates by $59\%$ compared to PaletteNeRF with semantic guidance. We use Segment Anything [18] to obtain masks for a subset of test set views and compare against PaletteNeRF [20]. We measure MSE compared to the ground truth test images and use the same hyperparameters as PaletteNeRF, which also builds on torch-ngp [40]. As can be seen, LAENeRF after recoloring outperforms non-recolored PaletteNeRF, which we attribute to PaletteNeRF’s concurrent scene reconstruction and palette-based decomposition. Due to the fairly accurate geometry for the $360^{\circ}$ captures, our approach introduces very few background artefacts during recoloring.
@@
In Fig. 4, we compare our method to ICE-NeRF [23] and PaletteNeRF [20]. As can be seen, our method introduces the fewest artefacts due to recoloring. To facilitate cross-method comparisons, we choose the same example as ICE-NeRF and include the results from their publication.
For the mip-NeRF 360 dataset [3], we compare against PaletteNeRF in Fig. 5. PaletteNeRF either introduces significant artefacts (see Room) or is unable to capture small regions with their semantic features (see Bonsai). As mentioned in their publication, ICE-NeRF struggles with local edits for this dataset, frequently introducing artefacts in undesired regions.
@@ 5.2 Non-Photorealistic Appearance Editing
For our quantitative evaluation, we measure MSE in the background with respect to the ground truth test set images. For NeRF-Synthetic [28], we generate masks for the region to stylize using our method and use segmentation masks from ICE-NeRF [23] for LLFF [28] scenes. We report per-dataset results in Tab. 3. In contrast to Ref-NPR, our approach demonstrates considerably reduced error rates, especially in synthetic scenes characterized by numerous occlusions.
For forward-facing scenes, Ref-NPR benefits from less variation between camera poses but still generates about 3×3$\times$3 × more errors than ours.
@@
Figure 8: User study results. Participants prefer our method to related work for image and video outputs. Our method is also rated higher for view-consistency than Ref-NPR [54]. $*{}^{*}$ indicates a statistical significance according to Wilcoxon signed rank tests.
@@ 7 Conclusion
We present LAENeRF, a unified framework for photorealistic and non-photorealistic appearance editing of NeRF. By elegantly combining a palette-based decomposition with perceptual losses, we enable interactive recoloring of stylized regions. We demonstrate state-of-the-art local appearance editing results, benefitting from our geometry-aware stylization in 3D. LAENeRF outperforms existing works quantitatively and qualitatively for local recoloring and local stylization. By open-sourcing our approach we will bring NeRF-editing to a large audience.
@@ A.3 Geometry-Aware Stylization
To perform stylization which respects the learned geometry of our pre-trained NeRF, we perform depth estimation. As we set the depth to 00 for all rays which did not intersect $\bm{\mathcal{E}}$, we would get incorrect results when computing our depth guidance image. Additionally, direct neighbors of rays which did not intersect the edit grid often do not accumulate full alpha inside $\bm{\mathcal{E}}$, leading to incorrect predictions for $0pt$. To remedy the aforementioned issues, we multiply ${\nabla 0pt}$ with the accumulated alpha inside the edit grid ${\alpha}^{{(\bm{\mathcal{E}})}}$, for both pixels involved in the computation and their direct neighbors:
[…]
@@ A.4 Modifications to our Framework
We include the following modifications to torch-ngp [40] to improve reconstruction for real-world scenes, following PaletteNeRF [20]. By default, the background color is assumed to be white for real-world scenes. We use a random background color during training, which leads to better depth estimation and a sparser occupancy grid. In addition, we mark grid cells between each camera and its’ corresponding near plane as non-trainable.
@@ Appendix C Additional User Study Details
We conduct our user study for a comparison with state-of-the-art methods for local recoloring and local style transfer in scenes represented by NeRFs. We conduct the study on an iPad 9th Gen with a 10.2” display. All images and videos used for our user study are included in the supplementary material.
@@
We compute the Wilcoxon signed rank test [48] to determine whether there is statistical significance in the preference for one or the other method. All preference scores indicate a significant preference for LAENeRF with recoloring ($Z=8850.0;\ p<0.0001$), stylization ($Z=5000.0;\ p<0.0001$), video recoloring ($Z=1872.5;\ p<0.0005$), and video stylization ($Z=1750.0;\ p<0.005$). The view consistency scores indicate a preference for our method for stylization ($Z=1232.0;\ p<0.001$), but no significant difference for recoloring ($Z=1890.0;\ p>0.5$).
@@ Appendix D GUI
LAENeRF incorporates a GUI building on [16, 40] for real-time, interactive appearance editing of NeRFs. Here, we present several key components of our user interface in detail.
@@ D.1 Region Selection and Growing
We provide an intuitive region selection process, based on NeRFShop [16]. First, the user scribbles on the image rendered from the current camera pose in our real-time viewer. For each selected ray, we perform raymarching and map the estimated ray termination to a the nearest voxel, which we set in $\bm{\mathcal{E}}$, representing our initial selection. Next, the initial selection can be extended with region growing, where we control growing by the number of voxels which we pop from our growing queue per-iteration and the total number of iterations. Our GUI shows the selection after each region growing step. We can optionally create a growing grid ${\bm{\mathcal{G}}}$ from our selection to model smooth transitions. Further, we incorporate binary operations with a second grid for more intuitive addition and deletion of selected voxels. The complete workflow can be seen in Fig. 14.
@@ D.3 More Palette Control
To provide users with more control over the stylized/recolored region, we enable a linear transformation of the learned weights ${\hat{{\bm{w}}}}$. To this end, we introduce palet weights ${{\bm{w}}_{{\hat{\bm{\mathcal{P}}}}}}\in\mathbb{R}^{{N_{{\hat{\bm{\mathcal{P}}}}}}}$ and palette biases ${{\bm{b}}_{{\hat{\bm{\mathcal{P}}}}}}\in[-1,1]^{{N_{{\hat{\bm{\mathcal{P}}}}}}}$, which transform the weights according to
@@ D.4 Preview Mode & GUI demonstration
To enable interactive recoloring and stylization, our approach supports real-time rendering during training of our LAENeRF module. After the training dataset has been extracted, which takes $\sim 15$ seconds depending on the dataset, users can watch our neural module converge. We prepare a video demonstration of our application, which is contained in the supplementary material. This demonstration also includes more recoloring and stylization results. The video is available here: https://youtu.be/iXVr6OCD_CA.

### v2

@@ 5.1 Photorealistic Appearance Editing
For the quantitative evaluation, we follow ICE-NeRF [23] and measure the Mean Squared Error (MSE) in the background of the selected region before and after recoloring. In Tab. 1, we present this metric for three scenes of the LLFF dataset [27]. The foreground region is described by a mask, which was provided to us by the authors of ICE-NeRF. We perform 7 different recolorings per scene and compare against PaletteNeRF with and without semantic guidance, whereas we also include the numbers from ICE-NeRF [^1] ICE-NeRFs’ implementation is not yet publicly available. to facilitate cross-method comparisons. LAENeRF outperforms previous methods for this metric, reducing error rates by $59\%$ compared to PaletteNeRF with semantic guidance. Additionally, we provide the same metric for the indoor scenes of the mip-NeRF 360 dataset [3] in Tab. 2, where we report average results for 7 recolorings per scene. We use Segment Anything [18] to obtain masks for a subset of test set views and compare against PaletteNeRF [20]. We measure MSE compared to the ground truth test images and use the same hyperparameters as PaletteNeRF, which also builds on torch-ngp [40]. As can be seen, LAENeRF after recoloring outperforms non-recolored PaletteNeRF, which we attribute to PaletteNeRF’s concurrent scene reconstruction and palette-based decomposition. Due to the fairly accurate geometry for the $360^{\circ}$ captures, our approach introduces very few background artefacts during recoloring.
@@
Zoomed-In Results
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
In Fig. 4, we compare our method to ICE-NeRF [23] and PaletteNeRF [20]. As can be seen, our method introduces the fewest artefacts due to recoloring. To facilitate cross-method comparisons, we choose the same example as ICE-NeRF and include the results from their publication. For the mip-NeRF 360 dataset [3], we compare against PaletteNeRF in Fig. 5. PaletteNeRF either introduces significant artefacts (see Room) or is unable to capture small regions with their semantic features (see Bonsai). As mentioned in their publication, ICE-NeRF struggles with local edits for this dataset, frequently introducing artefacts in undesired regions.
@@ 5.2 Non-Photorealistic Appearance Editing
For our quantitative evaluation, we measure MSE in the background with respect to the ground truth test set images. For NeRF-Synthetic [28], we generate masks for the region to stylize using our method and use segmentation masks from ICE-NeRF [23] for LLFF [28] scenes. We report per-dataset results in Tab. 3. In contrast to Ref-NPR, our approach demonstrates considerably reduced error rates, especially in synthetic scenes characterized by numerous occlusions. For forward-facing scenes, Ref-NPR benefits from less variation between camera poses but still generates about 3×3$\times$3 × more errors than ours.
@@
Figure 8: User study results. Participants prefer our method to related work for image and video outputs. Our method is also rated higher for view-consistency than Ref-NPR [53]. indicates a statistical significance according to Wilcoxon signed rank tests.
@@ 7 Conclusion
We present LAENeRF, a unified framework for photorealistic and non-photorealistic appearance editing of NeRF. By elegantly combining a palette-based decomposition with perceptual losses, we enable interactive recoloring of stylized regions. We demonstrate state-of-the-art local appearance editing results, benefiting from our geometry-aware stylization in 3D. LAENeRF outperforms existing works quantitatively and qualitatively for local recoloring and local stylization. By open-sourcing our approach we will bring NeRF-editing to a large audience.
@@
In this supplementary document, we provide more implementation details and quantitative results. Additionally, we describe our GUI in detail and provide additional details for our user study.
@@ A.3 Geometry-Aware Stylization
To perform stylization which respects the learned geometry of our pre-trained NeRF, we perform depth estimation. As we set the depth to 00 for all rays which did not intersect $\bm{\mathcal{E}}$, we would get incorrect results when computing our depth guidance image. Additionally, direct neighbors of rays which did not intersect the edit grid often do not accumulate full alpha inside $\bm{\mathcal{E}}$, leading to incorrect predictions for $0pt$. To remedy the aforementioned issues, we multiply ${\nabla 0pt}$ with the accumulated alpha inside the edit grid ${T_{\alpha}}$, for both pixels involved in the computation and their direct neighbors:
$({\nabla 0pt})_{i,j}=({\nabla 0pt})_{i,j}\cdot\left(\prod_{y=j-1}^{j+2}({T_{\alpha}})_{i,y},\ \prod_{z=i-1}^{i+2}({T_{\alpha}})_{z,j}\right).$ (19)
@@ A.4 Modifications to our Framework
We include the following modifications to torch-ngp [40] to improve reconstruction for real-world scenes, following PaletteNeRF [20]. By default, the background color is assumed to be white for real-world scenes — iNGP might cheat with this static background color by not accumulating full $\alpha$ along a ray. We use a random background color during training, which leads to better depth estimates and a sparser occupancy grid. In addition, we mark grid cells between each camera and its corresponding near plane as non-trainable.
@@ Appendix C Additional User Study Details
We conduct our user study for a comparison with state-of-the-art methods for local recoloring and local style transfer in scenes represented by NeRFs. We conduct the study on an iPad 9th Gen with a 10.2" display. All images and videos used for our user study are included in the supplementary material.
@@
We compute the Wilcoxon signed rank test to determine whether there is statistical significance in the preference for one or the other method. All preference scores indicate a significant preference for LAENeRF with recoloring ($Z=8850.0;\ p<0.0001$), stylization ($Z=5000.0;\ p<0.0001$), video recoloring ($Z=1872.5;\ p<0.0005$), and video stylization ($Z=1750.0;\ p<0.005$). The view consistency scores indicate a preference for our method for stylization ($Z=1232.0;\ p<0.001$), but no significant difference for recoloring ($Z=1890.0;\ p>0.5$).
@@ Appendix D GUI
LAENeRF incorporates a GUI building on [16, 40] for real-time, interactive appearance editing of NeRFs. Here, we present several key components of our user interface.
@@ D.1 Region Selection and Growing
We provide an intuitive region selection process, based on NeRFShop [16]. First, the user scribbles on the image rendered from the current camera pose in our real-time viewer. For each selected ray, we perform raymarching and map the estimated ray termination to a the nearest voxel, which we set in $\bm{\mathcal{E}}$, representing our initial selection. Next, the initial selection can be extended with region growing, where the user controls growing by the number of voxels which we pop from our growing queue per-iteration and the total number of iterations. Our GUI shows the selection after each region growing step. We can optionally create a growing grid ${\bm{\mathcal{G}}}$ from our selection to model smooth transitions. Further, we incorporate binary operations with a second grid for more intuitive addition and deletion of selected voxels. Due to the region growing procedure and easy deletion of voxels with a second grid and binary operations, our method is quite insensitive to the user scribble provided as input. The complete workflow can be seen in Fig. 14.
@@ D.3 More Palette Control
To provide users with more control over the stylized/recolored region, we enable a linear transformation of the learned weights ${\hat{{\bm{w}}}}$. To this end, we introduce palette weights ${{\bm{w}}_{{\hat{\bm{\mathcal{P}}}}}}\in\mathbb{R}^{{N_{{\hat{\bm{\mathcal{P}}}}}}}$ and palette biases ${{\bm{b}}_{{\hat{\bm{\mathcal{P}}}}}}\in[-1,1]^{{N_{{\hat{\bm{\mathcal{P}}}}}}}$, which transform the weights according to
@@ D.4 Preview Mode & GUI demonstration
To enable interactive recoloring and stylization, our approach supports real-time rendering during training of our LAENeRF module. After the training dataset has been extracted, which takes $\sim 15$ seconds depending on the dataset, users can watch our neural module converge. As we only handle views of the training dataset during optimization, this efficient dataset extraction results in fast convergence.
To render novel views during training, enabling interactive previews, we query $\Theta_{\text{NeRF}}$ to obtain the estimated ray termination ${\bm{x}}_{\text{term}}$ and ${T_{\alpha}}$, as we do during training dataset extraction. If ${T_{\alpha}}>0.5$, we use ${\bm{x}}_{\text{term}}$ as input to LAENeRF and output the rendered image. Consequently, due to the fast stylization process, users have the ability to stop training after few seconds and choose another style image if they find the result unappealing.